# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v2

@@ V-A3 Differential Privacy Techniques
In data analysis and machine learning, DP offers a robust framework for safeguarding individual privacy by introducing carefully calibrated noise into data or computations [181]. This noise injection allows meaningful insights to be extracted while preserving privacy, making DP particularly relevant in edge-cloud collaborative AI systems. Specifically, DP mechanisms ensure negligible impact on analysis outputs from the presence or absence of a single individual’s data, preventing identification or inference of sensitive information [182, 183]. Common approaches include adding noise directly to raw data before analysis or perturbing the results of computations, such as model parameters, with the noise calibrated based on the computation’s sensitivity [182, 183]. For instance, in federated learning, DP protects individual device data during collaborative training by adding noise to local model updates before aggregation, preventing private data reconstruction from shared updates [184]. Similarly, in collaborative inference, DP can protect intermediate outputs and mitigate membership information leakage [185]. While noise addition inevitably impacts data utility and model accuracy [186], a key challenge is balancing privacy and utility, especially given the heterogeneous nature of edge devices and distributed edge-cloud systems [187, 188].
@@ V-A4 Data Anonymization Approaches
In AI-driven ECCC environments, data anonymization techniques are crucial for maintaining privacy by transforming data to prevent individual identification [189]. Such environments, characterized by distributed data processing and model training, present unique privacy challenges, necessitating anonymization to safeguard sensitive information while enabling collaborative AI development. Several approaches achieve this goal, including StyleID, which anonymizes faces in image datasets through identity disentanglement [190]. Specifically, StyleID projects images onto a GAN latent space, manipulating specific features to protect identities while preserving other image characteristics. Another key technique, DP, adds noise to data or computations to protect individual privacy while preserving useful insights [86]. For instance, differentially private federated tensor completion enables secure cloud-edge collaborative AIoT data prediction [86]. Dynamic anonymization techniques, often leveraging machine learning, adapt to specific contexts and data characteristics for more flexible and effective privacy protection [189]. Data anonymization in edge-cloud collaborative AI systems, where data may be distributed across multiple devices and locations, can be combined with other privacy-preserving techniques. These techniques include SMPC and federated learning to create a more comprehensive privacy framework [86]. As an example, in cloud-edge collaborative inference, anonymization complements methods like PrivaScissors, which reduces mutual information leakage between intermediate model outputs and device data [185].
@@ V-B Model Security
TABLE VI: Summary of model integrity protection, attack mitigation, and secure update mechanisms.
Method Type Domain Performance Metrics Edge Cloud ECC Privacy Security
Li et al. [191] Data Auditing Next-Generation Consumer Electronics Data Integrity, Privacy
Alaverdyan et al. [192] Security Framework Edge-Cloud Networks Security, Data Protection
Chi et al. [193] Trusted Architecture AIoT Model Performance, Security
FlexibleFL [88] Defense Framework Cloud-Edge Federated Learning Defense Performance, Model Quality
Yao et al. [194] Data Sharing Terminal-Edge-Cloud Network Data Security, Access Control, Integrity
Yang et al. [86] Privacy Framework Cloud-Edge AIoT Data Prediction Accuracy, Privacy
Joshi et al. [195] Edge-AI Framework Landslide Monitoring Response Latency, Energy Consumption, Prediction Accuracy
Khuen et al. [80] Federated Learning Precision Aquaculture Bandwidth Usage, Data Privacy
Nguyen et al. [196] Resource Allocation Space-Air-Ground Integrated Networks (SAGIN) Energy Consumption, Task Delay
1D CNN-LSTM [197] Machine Learning Smart Education Systems Learning Effectiveness, Content Personalization
Shi et al. [198] Resource Management IIoT Energy Consumption, Processing Accuracy, Service Delay
Zeng et al. [199] QoS Testing Satellite Edge Clouds QoS Metrics for Microservices
RLDA [200] Data Aggregation Edge-Cloud IoT Systems Overhead with Growing Dropped Users, Security
Al-Hawawreh et al. [201] Security Framework IIoT Attack Detection Efficiency
Chen et al. [202] Security Framework Distributed Cyber-Physical Systems Attack Detection Efficiency, Resource Utilization
Fan et al. [203] Intrusion Detection Vehicular Networks Attack Detection Accuracy
GS-SNC [204] Secure Data Transmission AIoT Security, Computational Complexity, Encryption Capacity
Ke et al. [205] Reputation Assessment Edge VANETs Precision, Recall
Li et al. [206] Attack Detection Distribution Networks Detection Accuracy
ERLG [207] Game Theory Sensor Edge Cloud Alarm Message Reduction
Qiu et al. [208] Service Chain Provisioning Edge Intelligence Cloud Throughput, Deployment Cost
CD-DOACIR [209] Access Control Cloud-Edge IoT Effectiveness, Efficiency
Yu et al. [210] Game Theory 5G Distributed Networks Defense Effectiveness, Resource Allocation
KubeEdge-Sedna [211] Lifelong Learning Edge AI Adaptability to Heterogeneous Data
Liu et al. [212] Blockchain Framework Edge AIGC Product Protection, Trustworthiness
Abdelmoniem [213] Model-Centric Design Edge AI/ML Scalability, Collaboration Efficiency
TMP [214] Confidential Inference Cloud AI Service Information Leakage Prevention, Performance
Tidjon and Khomh [215] Framework AI-based Systems Trust Properties, Verification
Hutiri and Ding [216] Unified Framework Voice-activated Edge Services Privacy, Security, Fairness
CHAI-T [217] Process Framework Human-AI Collaboration Trust Evolution, Team Processes
U-Federator [76] Authentication Cloud, Edge, and Fog Computing Authentication Time
Hutiri et al. [216] Framework Edge Intelligence, Voice-activated Services Privacy, Security, Fairness
Distributed and collaborative nature of edge-cloud systems introduces significant security vulnerabilities for the AI models deployed within them, potentially exposing them to unauthorized modifications, adversarial manipulations, or inference attacks. Consequently, establishing robust security measures throughout the model lifecycle is paramount for ensuring trustworthy and reliable AI operations in ECCC environments. We explore critical security challenges through three key dimensions: model integrity, adversarial attack mitigation, and secure model updates. Initially, Sec. V-B1 investigates techniques to protect model integrity from tampering and corruption. Following this, Sec. V-B2 analyzes methods to detect and counter adversarial attacks and model poisoning that aim to degrade performance or manipulate behavior. Concluding the discussion, Sec. V-B3 focuses on secure protocols for model updates across the edge-cloud continuum, ensuring authenticity and minimizing vulnerabilities. To further illustrate these security mechanisms, Table VI compares recent approaches in ECCC systems. Notably, the table reveals a wide range of domains—from AIoT to vehicular networks—alongside consistent emphasis on privacy and security, with methods like FlexibleFL [88] demonstrating robust defense performance in federated learning scenarios.
V-B1 Model Integrity Protection
Model integrity protection is crucial in edge-cloud collaborative AI systems to ensure that models are not tampered with or corrupted, either accidentally or maliciously [88]. This protection encompasses various methods, focusing on both the storage and the usage of the AI models. Secure model storage mechanisms aim to safeguard models from unauthorized access and modification while at rest. Cryptographic hashing, for instance, can be employed to generate a unique fingerprint of the model [191]. Any subsequent alteration to the model will result in a different hash value, thereby enabling the detection of tampering. Furthermore, secure enclaves and trusted execution environments (TEEs) offer hardware-based isolation and protection for model storage and execution, preventing unauthorized access and modification [192]. Beyond secure storage, ensuring the integrity of models during operation is equally critical. Techniques such as code signing and model authentication can verify the origin and authenticity of models before they are loaded and executed [193]. Moreover, runtime monitoring can detect anomalies in model behavior, potentially indicating a compromise or corruption [194]. These methods, collectively, contribute to establishing a robust security framework for protecting the integrity of AI models in edge-cloud environments.
V-B2 Attack Detection and Prevention
AI models in edge-cloud systems are vulnerable to attacks, necessitating robust defense mechanisms. Adversarial attacks introduce subtle perturbations to input data, causing misclassification or incorrect outputs [203]. Detecting such attacks involves analyzing input data for anomalies; meanwhile, prevention strategies include adversarial training with adversarial examples and input sanitization to neutralize perturbations. Auto-updating mechanisms can further enhance adaptability to evolving attack strategies [203]. Model poisoning attacks, conversely, inject malicious data into training datasets, manipulating model learning and creating exploitable backdoors [88]. Detecting these attacks requires careful examination of training data for inconsistencies. Federated learning offers inherent protection by distributing training [86], and contribution-based mitigation strategies in frameworks like FlexibleFL [88] enhance resilience by evaluating uploaded model parameters. Secure network coding schemes [204] can protect data transmission, while trust management mechanisms [205] ensure model and update reliability.
V-B3 Secure Model Updates
Maintaining the integrity and trustworthiness of AI models in edge-cloud collaborative systems necessitates secure model updates [218]. Authentication mechanisms are essential for verifying the update source and preventing unauthorized modifications [219]. Specifically, digital signatures, cryptographic hashes, or blockchain-based solutions [212] can be employed for this purpose. Additionally, verification of update integrity ensures that updates remain untampered during transmission or storage [191], leveraging techniques such as checksums, message authentication codes (MACs), or Merkle trees. Secure transmission through encrypted channels using protocols like TLS or SSH [214] protects against eavesdropping. A trusted execution environment (TEE) [214] can further secure the update process itself from runtime attacks. Rollback mechanisms to a previous stable state in case of faulty updates are also crucial [211].
V-B4 Trust Management Mechanisms
In edge-cloud collaborative systems, ensuring the reliability and trustworthiness of AI models and their updates is crucial [220]. Specifically, this necessitates establishing mechanisms for evaluating and maintaining trust in the components and processes involved in model development, deployment, and operation. Given the distributed and collaborative nature of these environments, trust management must also address potential risks associated with interactions between edge devices, cloud servers, and the AI models themselves. Reputation systems are vital for establishing trust within collaborative AI environments [215]. They track and aggregate feedback on the performance and behavior of different entities, providing a basis for assessing trustworthiness. For example, an edge device’s reputation could be based on its history of providing accurate data and reliably executing tasks, while an AI model’s reputation could reflect its accuracy, fairness, and robustness [221]. These reputation scores subsequently inform decisions about trust allocation, resource distribution, and task assignment. Trust evaluation methods quantify and assess the trustworthiness of AI models and other system components [221]. Such methods can involve quantitative metrics like accuracy, latency, and energy efficiency, combined with qualitative assessments such as user feedback and expert reviews [217]. Context-specific application requirements also influence trust evaluation. For instance, safety-critical applications, like autonomous driving, prioritize robustness and reliability [222], whereas applications like personalized healthcare prioritize privacy and fairness [216]. Robust and context-aware trust evaluation is therefore crucial, including consideration of the trustworthiness of AI engineering schemes themselves, as exemplified by platforms like KubeEdge-Sedna [211].
@@ V-C Communication Security
Securing communication channels is paramount in edge-cloud collaborative systems to protect data transmission from unauthorized access and tampering. This protection involves secure data transmission methods, including encryption and secure communication protocols, safeguarding data in transit. Authentication and access control mechanisms, such as multi-factor authentication and role-based access control, prevent unauthorized access to resources and data. Various encryption schemes, including symmetric and asymmetric encryption, secure data storage and transmission in edge-cloud environments. Additionally, blockchain technology offers the potential to enhance security by providing decentralized security, data integrity, and transparent audit trails.
V-C1 Secure Data Transmission
Protecting sensitive information in transit between edge devices and cloud servers requires robust security mechanisms in ECCC systems [223]. Various encryption methods serve critical roles in securing data transmission, with symmetric encryption offering efficiency for large data transfers through a single key approach, while asymmetric encryption enhances security for key exchange by employing separate encryption and decryption keys [224, 225]. Secure communication protocols including transport layer security (TLS) and its predecessor secure sockets layer (SSL) establish protected channels that ensure data confidentiality and integrity by utilizing cryptographic algorithms to encrypt data during transmission [226]. Additional security layers can be implemented through virtual private networks (VPNs) that create secure connections over public networks [227], while blockchain technology provides decentralized and immutable frameworks for creating transparent audit trails that prevent unauthorized modifications and strengthen the security of edge-cloud communication [194, 192].
V-C2 Authentication and Access Control
In edge-cloud collaborative systems, robust security is paramount due to their distributed nature and the potential involvement of resource-constrained devices. Specifically, authentication and access control mechanisms are essential for preventing unauthorized access to data and resources [228]. Without these safeguards, malicious actors could compromise sensitive data or disrupt system operations, while inadequate access control could lead to data breaches or system instability due to authorized users accessing resources beyond their permitted scope. Authentication verifies user or device identity attempting system access, while access control governs permitted resources and actions [209]. This combined approach is crucial for maintaining security and integrity in these environments. For example, DOACIR [209] and P-MOD [229] offer enhanced security through controlled data sharing and hierarchical privilege management, respectively. Multi-factor authentication (MFA) strengthens security by requiring multiple forms of identification, mitigating risks even if one factor is compromised. Additionally, role-based access control (RBAC) [230] streamlines access management by assigning permissions to roles, improving efficiency and reducing errors, especially in complex systems. Integrating blockchain technology can further enhance RBAC by providing a secure platform for managing access control policies [230].
V-C3 Encryption Schemes
Robust encryption schemes are essential for protecting data both in transit and at rest within edge-cloud collaborative environments. These schemes are broadly classified as symmetric and asymmetric encryption, each possessing distinct advantages and applications. Symmetric encryption uses a single key for both encryption and decryption, offering high performance suitable for securing large data volumes, such as data streams between devices and edge nodes [194]. This efficiency is particularly beneficial for resource-constrained edge devices, although secure key distribution and management present challenges. In contrast, asymmetric, or public-key, encryption utilizes separate keys for encryption and decryption, simplifying key distribution and facilitating secure communication across the edge-cloud hierarchy [192]. Asymmetric encryption also enables fine-grained access control, as exemplified by CP-ABE schemes, where data decryption is contingent on possessing specific attributes [209]. While offering enhanced security, asymmetric encryption typically incurs higher computational overhead. Consequently, edge-cloud systems often combine symmetric and asymmetric encryption to leverage the strengths of both, potentially integrating hybrid models with blockchain technology for enhanced data integrity and auditability in cloud storage [194].
V-C4 Blockchain-based Security Solutions
We explore how blockchain technology enhances security in edge-cloud collaborative environments through decentralized architecture and immutable data records. In edge-cloud collaborative systems, blockchain offers a promising approach to enhancing security [231]. Its decentralized nature, based on cryptographic principles, distributes security responsibility across a network of nodes, eliminating the single point of failure vulnerability inherent in centralized systems [232]. Consequently, this distributed architecture enhances resilience against attacks and ensures system availability. Blockchain’s immutability guarantees data integrity, preventing unauthorized modifications and ensuring data provenance [233, 234]. Specifically, once a transaction is recorded on the blockchain, it becomes tamper-proof. Additionally, blockchain facilitates transparent audit trails by recording all transactions on a public ledger [235]. This characteristic enables verifiable tracking of data flow and system operations, promoting accountability and facilitating efficient auditing for security and compliance [194].
@@ VI-B Smart Healthcare
Figure 7: Integration of federated learning, edge-cloud platforms, and Internet of medical things (IoMT) optimization in smart healthcare.
Advancements in AI and edge-cloud computing are transforming smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig.7. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [241]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [242]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [243], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, process it at the edge for immediate feedback, and transmit it to the cloud for analysis and storage [25]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [244]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [245]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [246]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [247].
@@ VI-C Industrial Automation
The integration of ECCC delivers significant advantages to industrial automation processes. Edge intelligence enables real-time fault diagnosis and predictive maintenance [18], while anomaly detection techniques deployed on AIoT devices enhance system reliability by identifying unusual patterns and potential failures before they cascade throughout the system [246, 16]. Leveraging both cloud computational power and edge real-time processing capabilities, industrial automation systems achieve greater agility through optimized production processes and resource allocation [70], where cloud-based platforms analyze historical data to generate optimized production schedules while edge devices adapt these schedules in real-time based on local conditions. Distributed intelligence across multiple edge devices handles complex manufacturing tasks more efficiently and enhances fault tolerance since individual device failures do not necessarily compromise the entire system [248]. As edge intelligence, anomaly detection, cloud-edge collaboration, and distributed intelligence converge, industrial automation systems evolve toward more sophisticated and autonomous manufacturing processes [56].
@@ VI-D Smart Cities
ECCC enhances urban services in smart cities by enabling intelligent traffic management and environmental monitoring [249, 133]. Edge devices collect and analyze real-time data for adaptive traffic control and congestion reduction, while distributed sensor networks monitor environmental factors like air quality and noise levels. Subsequently, cloud platforms aggregate and analyze this data to inform urban planning and environmental protection. In addition, AI-driven resource allocation optimizes smart grid energy consumption [250]. Specifically, edge devices monitor energy usage, and cloud-based algorithms dynamically adjust energy distribution to minimize waste and improve grid stability. FL is crucial for privacy-preserving data analysis and decision-making in these applications [13, 251, 252]. Through FL, stakeholders such as government agencies and private companies can collaboratively train models on combined data without sharing sensitive information, leading to more accurate and robust models for urban planning and public safety.

### v4

@@ V-A3 Differential Privacy Techniques
Privacy-preserving analysis frameworks offer robust protection for individual privacy by introducing carefully calibrated noise into data or computations [178], enabling meaningful insights extraction while safeguarding sensitive information. Specifically, DP mechanisms ensure negligible impact on analysis outputs from the presence or absence of a single individual’s data, preventing identification or inference of sensitive information [179, 180]. Common approaches include adding noise directly to raw data before analysis or perturbing the results of computations, such as model parameters, with the noise calibrated based on the computation’s sensitivity [179, 180]. For instance, in federated learning, DP protects individual device data during collaborative training by adding noise to local model updates before aggregation, preventing private data reconstruction from shared updates [181]. Similarly, in collaborative inference, DP can protect intermediate outputs and mitigate membership information leakage [182]. While noise addition inevitably impacts data utility and model accuracy [183], a key challenge is balancing privacy and utility, especially given the heterogeneous nature of edge devices and distributed edge-cloud systems [184, 185].
@@ V-A4 Data Anonymization Approaches
Identity protection techniques transform data to prevent individual identification while enabling collaborative AI development, particularly crucial in distributed processing environments with unique privacy challenges [186]. Several approaches achieve this goal, including StyleID, which anonymizes faces in image datasets through identity disentanglement [187]. Specifically, StyleID projects images onto a GAN latent space, manipulating specific features to protect identities while preserving other image characteristics. Another key technique, DP, adds noise to data or computations to protect individual privacy while preserving useful insights [97]. For instance, differentially private federated tensor completion enables secure cloud-edge collaborative AIoT data prediction [97]. Dynamic anonymization techniques, often leveraging machine learning, adapt to specific contexts and data characteristics for more flexible and effective privacy protection [186]. Data anonymization in edge-cloud collaborative AI systems, where data may be distributed across multiple devices and locations, can be combined with other privacy-preserving techniques. Complementary approaches include SMPC and federated learning to create a more comprehensive privacy framework [97]. As an example, in cloud-edge collaborative inference, anonymization complements methods like PrivaScissors, which reduces mutual information leakage between intermediate model outputs and device data [182].
@@ V-B Model Security
Comprehensive security frameworks address critical vulnerabilities throughout AI model lifecycles via multi-layered protection mechanisms against unauthorized modifications, adversarial attacks, inference attacks, and model poisoning [188, 97]. Robust protection encompasses integrity mechanisms utilizing cryptographic hashing and trusted execution environments, attack detection through adversarial training and anomaly detection, and secure update protocols implementing authentication and encrypted transmission [189, 190, 191]. Advanced approaches including FlexibleFL demonstrate effective federated learning defense through contribution-based mitigation, while blockchain-based solutions provide enhanced protection against coordinated attacks [99, 192, 193]. Additionally, runtime monitoring and auto-updating mechanisms enable adaptive responses to emerging threats while maintaining system resilience against novel attack patterns [194, 195]. Detailed technical analysis of integrity protection, attack detection, secure updates, and trust management is provided in LABEL:sec:supp-model-security of the Supplementary Material.
@@ V-C Communication Security
Secure data exchange across heterogeneous infrastructures requires comprehensive protection through integrated encryption, authentication, and access control mechanisms addressing distributed edge-cloud architectural challenges [196, 197]. Adaptive security strategies balance protection strength with computational efficiency through frameworks integrating symmetric encryption for large-scale transfers, asymmetric encryption for secure key exchange, and TLS/VPN protocols establishing protected channels [198, 199]. Authentication incorporates multi-factor authentication, role-based access control, and hierarchical privilege management preventing unauthorized access while maintaining operational efficiency [200, 201]. Furthermore, encryption schemes leverage attribute-based approaches for fine-grained access control, while blockchain-based solutions provide decentralized trust management and immutable audit trails [194, 202, 203]. Coverage of secure transmission protocols, authentication mechanisms, encryption schemes, and blockchain solutions is detailed in LABEL:sec:supp-communication-security of the Supplementary Material.
@@ VI Applications
ECCC enables transformative capabilities across diverse application domains by leveraging the unique synergy between edge proximity and cloud resources. Unlike traditional centralized or purely edge-based approaches, ECCC provides distinctive advantages through: (1) Hierarchical intelligence deployment, sophisticated AI models in the cloud guide lightweight edge models for real-time decision-making; (2) Dynamic workload distribution, computational tasks are intelligently partitioned based on latency requirements, privacy constraints, and resource availability; (3) Collaborative learning, federated and continual learning approaches enable privacy-preserving model improvement across distributed deployments; (4) Adaptive resource orchestration, dynamic scaling between edge and cloud resources based on real-time demands and network conditions. The following subsections demonstrate how these ECCC-specific capabilities create unprecedented opportunities for innovation across autonomous driving, smart healthcare, industrial automation, and smart cities, highlighting applications that are either impossible or significantly limited without the collaborative edge-cloud paradigm.
@@ VI-B Smart Healthcare
Figure 6: Integration of federated learning, edge-cloud platforms, and IoMT optimization in smart healthcare.
AI-driven collaborative computing transforms smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig. 6. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [209]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [210]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [211], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, process it at the edge for immediate feedback, and transmit it to the cloud for analysis and storage [28]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [212]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [213]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [214]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [215].
@@ VI-C Industrial Automation
Intelligent automation systems benefit significantly from collaborative computing architectures that deliver real-time fault diagnosis and predictive maintenance capabilities [21]. Leveraging both cloud computational power and edge real-time processing capabilities, industrial automation systems achieve greater agility through optimized production processes and resource allocation [80], where cloud-based platforms analyze historical data to generate optimized production schedules while edge devices adapt these schedules in real-time based on local conditions. Distributed intelligence across multiple edge devices handles complex manufacturing tasks more efficiently and enhances fault tolerance since individual device failures do not necessarily compromise the entire system [216]. As edge intelligence, anomaly detection, cloud-edge collaboration, and distributed intelligence converge, industrial automation systems evolve toward more sophisticated and autonomous manufacturing processes [65].
@@ VI-D Smart Cities
Urban infrastructure optimization leverages collaborative computing to enhance traffic management and environmental monitoring capabilities [217, 142]. Edge devices collect and analyze real-time data for adaptive traffic control and congestion reduction, while distributed sensor networks monitor environmental factors like air quality and noise levels. Subsequently, cloud platforms aggregate and analyze this data to inform urban planning and environmental protection. In addition, AI-driven resource allocation optimizes smart grid energy consumption [218]. Specifically, edge devices monitor energy usage, and cloud-based algorithms dynamically adjust energy distribution to minimize waste and improve grid stability. FL is crucial for privacy-preserving data analysis and decision-making in these applications [16, 219, 220]. Through FL, stakeholders such as government agencies and private companies can collaboratively train models on combined data without sharing sensitive information, leading to more accurate and robust models for urban planning and public safety.